Assume that a manufacturer can purchase a needed component from a supplier at a cost of $8 per unit, or it can invest $40,000 in equipment and produce the item at the cost of $5.50 per unit.
i) Determine the quantity for which total costs are equal for the make and buy alternatives.
ii) What is the minimum cost alternative if 15,000 units are required? What is the minimum cost?

PLEASE HELP ME FIND IT, THANKS.

Answer by jorel555(1290)

i) 8x=5.50x+40000
2.5x=40000
x=40000/2.5=16000 units, at which total costs are equal.
ii)8x15000 units=$120000
(15000x5.5)+40000=$122500 cost making the component oneself. The minimum cost is $120000 when buying the pre-made component, obviously!!!!!!!!!